1. The average number of milligrams(mg) of cholesterol in a cup of a certain brand of ice cream is 660 Ng, the standard deviation is 35mg. Assume the variable is normally distributed.a.If a cup of ice cream is selected, what is the probability that the cholesterol content will be more than 670mg?


1
Expert's answer
2021-05-24T18:07:48-0400

"P(X>670)=P(Z>\frac{670-660}{35})=P(Z>0.29)=1-P(Z<0.29)="

"=1-0.6141=0.3859."